Solve the equation by making an appropriate substitution.
x^2/3-4x^1/3+3=0

Answer by stanbon(75887)

Solve the equation by making an appropriate substitution.
x^2/3-4x^1/3+3=0
------
Let w = x^(1/3)
------
Substitute to get:
w^2-4w+3 = 0
(w-3)(w-1) = 0
w = 3 or w = 1
----
Solve for "x":
x^(1/3) = 3
x = 27
----
x^(1/3) = 1
x = 1
----------
Cheers,
Stan H.
===========